OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers ? A . 18 B. 19 C. 20 D. 21 ?

OpenStudy (hoblos):

x+y=51 x-y=13 find y

OpenStudy (anonymous):

\[x+y=51\] \[x-y=13\] add to get \[2x=64\] \[x=32\]

OpenStudy (hoblos):

yeah x=32 then y=32-13=19

OpenStudy (anonymous):

x+y=51 x-y=13 =>2y=51-13 =>y=38/2 =>y=19

OpenStudy (ragingsquirrel):

TL:DR B is the answer